I am having lots of trouble answering this equation.
Solve by substitution method:
4x + 3y = 10
y = 2x
Can you please help me?
Thank you!

Answer by Earlsdon(6103)

Solve by substitution:
4x+3y = 10
y = 2x Substitute this for the y in the 1st equation and solve for x.
4x+3(2x) = 10 Simplify.
4x+6x = 10 Combine like-terms.
10x = 10 Divide by 10.
x = 1 Now substitute this for the x in the 2nd equation.
y = 2(1)
y = 2
Solution is: (1, 2)